### v3

@@ 1 Introduction
In certain cases, it is logical to assume that the reference is observable. For instance, for a DM in the market for a new car, their current car can be a reference. However, in many situations, it is not realistic to assume the knowledge of the DM’s reference to anyone else but the DM themselves. Further, to analyze a choice problem with observable references, we have to deal with a family of stochastic choice functions, as done by Kovach and Suleymanov (2023). The inputs for the stochastic choice functions in their model were pairs of subsets [^3] Hereafter referred to as menus. and references- one family of stochastic choice function for each alternative as the reference. This is quite a significant data requirement. If we restrict the references to be unknown, we deal with just a single stochastic choice function over menus and thus, we do not have to deal with families of stochastic choice functions for each possible reference.
@@
It is also not obvious that a reference alternative must be unique. In what follows, I allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged, i.e. it is always paid attention to. Since I am looking at choices from menus that depend on unobservable reference alternatives, I assume that the references are menu-independent. It is certainly plausible that references can be menu-dependent, through which we can explain a larger set of stochastic choice behavior, but dealing with menu-dependent unobservable references is beyond the scope of this paper. I provide below three examples that match the mentioned assumptions.
  • Example 1: Michael Jordan is associated with the Air Jordan brand. A fan of his who is interested in buying shoes might consider an Air Jordan shoe as a reference. An outside observer need not know that the Air Jordan is in fact a reference for the DM. It is also likely that it stays as a reference whenever the DM has to make a shoe choice, so it is menu-independent. Finally, it does not preclude the DM to also be a fan of another athlete which allows this DM to have multiple reference alternatives.
  • Example 2: Suppose your neighbor buys a new flat-screen 4K TV and invites you over to watch a football game together. You are impressed by the picture quality and ask about the product. The prior knowledge of this particular TV will mean that this product will act as a reference point the next time you buy a TV. Knowledge of such a reference to the outside observer is unlikely, and it also keeps the possibility of multiple references intact.
@@
  • Example 3444I would like to thank Marco Mariotti for this application of the model.: Suppose past choices of the DM influence their future choices, i.e. they act as a reference. The outside observer only observes the current feasible sets and the choices from those but not the choices in the past. In this case, (multiple) past choices act as references and are thus, unobservable to the researcher.
@@ 2 Random Attention Model with Unobserved References
Let $X$ be the finite set of all possible alternatives. Denote the set of all non-empty subsets of $X$ by $\mathcal{X}$, with a typical element being $A$.
@@
A stochastic choice function is a mapping $p:X\times\mathcal{X}\rightarrow[0,1]$ where for each $A\in\mathcal{X}$ the following three conditions hold:
@@
Below I denote by $a^{*}$ the default alternative and interpret it as abstention or failing to make a choice. The idea of default is common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^6] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $A\in\mathcal{X}$, $A^{*}=A\cup\{a^{*}\}$.
@@
A stochastic choice function $p$ is a Random Attention Model (RAM) if there exists an attention function $\mu:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ and a strict total order $\succ$ [^7] A strict total order is an asymmetric, transitive and complete binary relation defined over $X$. over $X$ such that for all $A\in\mathcal{X}$ and $x\in A$:
@@
The idea here is similar to attention privilege as defined by Kovach and Suleymanov (2023), i.e., if a reference alternative is in the menu, it is always considered. In fact, that is exactly how I define a reference alternative in this setting. A reference alternative can have other properties, which maybe correlated with the preferences, dependent upon some status quo, or some attribute (the color of a car). However, the presented model is silent on any such dependence. An alternative is a reference if and only if it is attention privileged, irrespective of how it attained such a privilege. In my model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by Kovach and Suleymanov (2023) since the primitives there are menu-reference pairs. Note that the default is only chosen in case none of the alternatives are paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$ and allows us to interpret the default as a reference.
@@
A stochastic choice function $p$ is a random attention model with unobserved references (RAM-UR) if there exists a set $E\subseteq X$, a strict total order $\succ$ over $X$ and a reference-dependent attention function $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that for all $A\in\mathcal{X}$ and $x\in A$:
@@ 3 Characterization
In the present section I characterize the RAM-UR by using three testable axioms. These axioms restrict how the choices behave across different menus.
Existence of a Dominant Alternative (EDA): For each $A\in\mathcal{X}$: $p(x,A)=0$ for some $x\in A^{*}$ implies $p(y,\{x,y\})=1$ for some $y\in A\setminus\{x\}$.
@@ 4 Random Attention Model with Unobserved References: Independent Random Attention
In this section I assume that the attention function in the RAM-UR model is of an independent random type in the sense that there exists an attention parameter $\gamma:X\rightarrow(0,1]$ such that the attention function takes the form: $\mu_{IRA}(D,A)=\prod_{x\in D\subseteq A}\gamma(x)\prod_{y\in A\setminus D}(1-\gamma(y))$ for all $A\in\mathcal{X}$. Here attention privilege manifests in the form of alternatives with the attention parameter to be one, which is the key difference between this form and the attention function of the IRA form introduced by Manzini and Mariotti (2014) (hereafter referred to as the IRA model). The IRA restricts the attention parameter to be strictly between zero and one while we do not impose such a restriction. This makes the IRA model to be nested in the RAM-UR-IRA.
@@
A stochastic choice function $p$ is a random attention model with unobservable references and independent random attention (RAM-UR-IRA) if there exists a strict total order $\succ$ over $X$ and an attention parameter $\gamma:X\rightarrow(0,1]$ such that for all $A\in\mathcal{X}$ and $x\in A$:
@@
Imposing the above additional structure significantly changes how the probabilities can be defined in the RAM-UR-IRA model. Even though it is a special case of RAM-UR and nests [^11] By nesting the IRA (rational) I mean that a special case of RAM-UR-IRA: the one where $E=\phi$ will also be an IRA (rational) model. both the IRA as well as classical rationality, it is mathematically a small departure from the IRA model: It allows $\gamma$ to take value one for some alternatives. In what follows, I provide a characterization of this model by using axioms similar to the characterizations of both the IRA and the RAM-UR and discuss the identification results.
@@ 4.1 Characterization
R-Asymmetry (R-ASYM): For all $A\in\mathcal{X}$ and $a,b\in A$ with $p(a,A)>0$, $p(b,A)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}\neq 1$ implies $\frac{p(b,A\setminus\{a\})}{p(b,A)}=1$.
@@
R-Independence (R-IND): (1) For all $A,B\in\mathcal{X}$ and $a,b\in A\cap B$ with $p(a,A)>0,p(a,B)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,B\setminus\{b\})}{p(a,B)}$.
(2) For all $A,B\in\mathcal{X}$ and $b\in A\cap B$ with $p(a^{*},A)>0,p(a^{*},B)>0$: $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{p(a^{*},B\setminus\{b\})}{p(a^{*},B)}$.
@@
Existence of a Dominant Alternative* (EDA*): For all $A\in\mathcal{X}$: $p(a^{*},A)=0$, implies $p(a^{*},\{x\})=0$ for some $x\in A$.
@@
Regularity (REG): For all $A,B\in\mathcal{X}$: $A\subset B$ implies $p(a,A)\geq p(a,B)$ for all $a\in A^{*}$.
@@ 5 Literature Overview
Models of stochastic choice ventured into modelling the variability observed in choices due to limited attention. Manzini and Mariotti (2014) was the seminal paper that filled this void. Since then there has been a substantial body of literature combining both limited attention and stochastic choice. A plethora of papers are closely related to Manzini and Mariotti (2014), for instance, Brady and Rehbeck (2016)’s feasibility model is a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility. Another one is Horan (2019) which looks at the IRA without a default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and models a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. They discussed the nestedness properties among these special cases. They were unable to conduct the nestedness analysis with respect to the classical stochastic choice models because in their model, the primitives were menu-reference pairs, which placed it outside the realm of stochastic choice models.
The idea of reference-dependent choice has also been the focus of several papers. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^13] Masatlioglu and Ok (2005) provides a few references for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. They model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They study the existence of a linear order such that the choice from each menu-reference pair is at least as good as the reference point, though not necessarily the maximum [^14] The best alternative according to the preference ordering.. The RAM-UR does have this non-deteriorating property since references are attention privileged, so in any menu, an alternative worse than the reference is never chosen.
@@
We have assumed menu-independence with respect to the references. In Kovach and Suleymanov (2023), the input to the problem is menu-reference pairs so the question of menu independence was moot. However, it is possible to allow for menu-dependent unobservable references as a possible extension to the model in the current paper. This extension obviously has bite, i.e., we cannot explain any stochastic choice function through such a model. For example, any stochastic choice function that allows for alternative $x$ to be chosen with certainty in a menu where $y$ is present and allows alternative $y$ to be chosen with certainty in a menu where $x$ is present cannot be a reference-dependent RAM even with menu dependent references.
@@
I take the set $\bar{E}=\{x\in X|p(x,\{x\})=1\}$ to be the revealed set of reference alternatives. If This set is empty EDA ensures that for any $A\in\mathcal{X}$, $p(a,A)>0$ for all $a\in A^{*}$. In this case, we can construct an attention function with $E=\bar{E}=\phi$ and an arbitrary preference. Note that with an empty set of references, the RAM-UR is just a RAM, so it suffices to show a RAM representation. For this, I take an arbitrary preference and construct an attention function as follows: For any alternative the probability that it is chosen is taken to be the sum of the value that the attention function takes on subsets where that alternative is the best. One way to do it is divide the probability of an arbitrary alternative $x$ in a menu $A$ to all sets in $\{D\subseteq A:x\succ y,\forall y\in D\setminus\{x\}\}$ equally ($\succ$ is the arbitrary preference). Note that since all alternatives (and default) are chosen with positive probability in all menus such a construction will give us a valid attention function. Given the constructed attention function, the observed choices are represented by a RAM with the attention function and the arbitrary preferences. Such a construction will be one possible RAM representation. We can have multiple representations if we change how we assign the probabilities. Thus, we have a representation when $\bar{E}$ is empty. For the rest of the proof I assume that $\bar{E}$ is non-empty.
@@
R-ASYM: If there exist $A\in\mathcal{X}$ and $x,y\in A$ with $p(x,A)>0$ and $p(y,A)>0$, then either none of $x$ and $y$ are references, in which case the axiom is true because of the functional form, as proved in Manzini and Mariotti (2014), or exactly one of them is a reference, and the other is preferred to it. In the latter case, the antecedent is true for the reference, since the probability of choice of the less preferred alternative (the reference) will increase if the other is removed, but removing the reference will not affect the choice probability of the better alternative.
R-IND: If there exists $A,B\in\mathcal{X}$, $a,b\in A\cap B$ with $p(a,A)>0$ and $p(a,B)>0$, the functional form is such that the R-IND will hold. Same is true if instead of $a$ we consider the default $a^{*}$. The proof is identical to the proof of the necessity of i-independence in Manzini and Mariotti (2014).
@@
EDA*: If $p(a^{*},A)=0$, for some $A\in\mathcal{X}$ then there exists at least one alternative $x\in A$ such that $\gamma(x)=1$. Then $p(a^{*},\{x\})=0$ follows.
@@
I take the set $\bar{E}=\{x\in X:p(x,A)=1\hskip 1.42262pt\text{for some}\hskip 1.42262ptA\in\mathcal{X}\}$ to be the revealed set of reference alternatives.
Again, this set could be empty. If this set is empty, then EDA* and NT ensure that $p(a,A)>0$ for all $A\in\mathcal{X}$ and $a\in A^{*}$. Now the R-ASYM and the R-IND will be equivalent to the i-ASYM and i-IND by Manzini and Mariotti (2014), which gives us the required equivalence with the IRA model [^16] An IRA model is a RAM-UR-IRA model with $E=\phi$.. For the remainder of the proof, I assume non-emptyness of $\bar{E}$.
@@
Reference-Independence of Irrelevant Alternatives (RIIA): For all $A\in\mathcal{X}$ and $a\in A$: $p(a,A)=0$ implies $p(b,A\setminus\{a\})=p(b,A)$ for all $b\in A\setminus\{a\}$.
Claim 1: If a stochastic choice function $p:X\times\mathcal{X}\rightarrow[0,1]$ satisfies REG, then it satisfies RIIA.
@@
Expansion (EXP): For all $A,B\in\mathcal{X}$, $x_{1}\in A,x_{2}\in B$: $p(x_{1},A)=p(x_{2},B)=1$ implies either $p(x_{1},A\cup B)=1$ or $p(x_{2},A\cup B)=1$.
Claim 2: If a stochastic choice function $p:X\times\mathcal{X}\rightarrow[0,1]$ satisfies R-ASYM, R-IND and REG, then it satisfies EXP.
@@
$A\cap\bar{E}=\phi$ for some $A\in\mathcal{X}$ implies $p(a,A)>0$ for all $a\in A$.

### v4

@@ 1 Introduction
In certain cases, it is logical to assume that the reference is observable. For instance, for a DM in the market for a new car, her current car can be a reference. However, in many situations, it is not realistic to assume the knowledge of the DM’s reference to anyone else but the DM themselves. Further, to analyze a choice problem with observable references, we have to deal with a family of stochastic choice functions, as done by Kovach and Suleymanov (2023). The inputs for the stochastic choice functions in their model were pairs of subsets [^3] Hereafter referred to as menus. and references- one family of stochastic choice function for each alternative as the reference. This is quite a significant data requirement. If we restrict the references to be unknown, we deal with just a single stochastic choice function over menus and thus, we do not have to deal with families of stochastic choice functions for each possible reference.
@@
It is also not obvious that a reference alternative must be unique. In what follows, I allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged, i.e. it is always paid attention to. Since I am looking at choices from menus that depend on unobservable reference alternatives, I assume that the references are menu-independent. It is certainly plausible that references can be menu-dependent, through which we can explain a larger set of stochastic choice behavior, but dealing with menu-dependent unobservable references is beyond the scope of this paper. I provide below four examples that match the mentioned assumptions.
  • Example 1: University Choice: High School students making a choice of university to continue their higher education in the UK filter out the large number of choices into smaller consideration sets. The universities are divided into old and new universities. This exogenous distinction impacts what universities are considered by students. For instance, Dawes and Brown (2005) found that minority ethnic groups had a greater proportion of old universities in their consideration set, possibly due to reputation concerns of these ethnic groups. The college choice of high school students in the US is also a pretty involved process. Rosen et al. (2008) showed that what major is offered as well as location were crucial factors in shortlisting colleges. Parental and guidance counselor advice also has a significant impact on the decision-making process. It is conceivable that colleges that are close by act as references. Dawes and Brown (2005) observed a similar phenomenon in the UK data as well. If the guidance of a counselor or a parent impacts which colleges act as references, it is not possible to observe these references.
  • Example 2: Apparel Shopping: While shopping for apparel, brand name is often crucial in drawing attention to particular items. The Air Jordan always catches the eye due to its association with Michael Jordan. However, though brand awareness might be enough to pull attention, it is not enough to provide a positive brand association as found by Dew and Kwon (2009). Thus, attention and preference are independent attributes.
  • Example 3: Political Choice: A stochastic choice framework can also be used to model a population of voters with similar preferences but with access to different options. A nationally strong party (a reference) will always be considered; however, a weaker party in one district might be a stronger party in another and be considered there, leading to different choices even when voters exhibit homogenous preferences. Wilson (2008) showed this for the Mexico national elections in 2000. The biggest party in Mexico, the Partido Revolucionario Institucional (PRI) has a strong presence in all the states, which makes it the reference, while the two opposition parties have varied presence in different states.
@@
  • Example $4$I would like to thank Marco Mariotti for this interpretation of the model.: Suppose past choices of the DM influence their future choices, i.e. they act as a reference. The outside observer only observes the current feasible sets and the choices from those but not the choices in the past. In this case, (multiple) past choices act as references and are thus, unobservable to the researcher. If the reference depends on recommendations from advisors (college choice) or past choices in elections (political choice), it is possible that the researcher is not able to observe the reference.
@@ 2 Random Attention Model with Unobserved References
Let $X$ be the finite set of all possible alternatives. Denote the set of all subsets of $X$ by $\mathcal{X}$, with a typical element being $A$.
@@
A stochastic choice function is a mapping $p:X\times\mathcal{X}\setminus\phi\rightarrow[0,1]$ where for each $A\in\mathcal{X}\setminus\phi$ the following three conditions hold:
@@
Below I denote by $a^{*}$ the default alternative and interpret it as abstention or failing to make a choice or delaying choice. The idea of default is common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^6] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}\setminus\phi$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $A\in\mathcal{X}\setminus\phi$, $A^{*}=A\cup\{a^{*}\}$.
@@
A stochastic choice function $p$ is a Random Attention Model (RAM) if there exists an attention function $\mu:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ and a strict total order $\succ$ [^7] A strict total order is an asymmetric, transitive and complete binary relation defined over $X$. over $X$ such that for all $A\in\mathcal{X}\setminus\phi$ and $x\in A$:
@@
The idea here is similar to attention privilege as defined by Kovach and Suleymanov (2023), i.e., if a reference alternative is in the menu, it is always considered. In fact, that is exactly how I define a reference alternative in this setting. A reference alternative can have other properties, which maybe correlated with the preferences, dependent upon some status quo, or some attribute (for instance, the color of a car). However, the presented model is silent on any such dependence. An alternative is a reference if and only if it is attention privileged, irrespective of how it attained such a privilege. In my model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by Kovach and Suleymanov (2023) since the primitives there are menu-reference pairs. Note that the default is only chosen in case none of the alternatives are paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$ and allows us to interpret the default as a reference.
@@
A stochastic choice function $p$ is a random attention model with unobserved references (RAM-UR) if there exists a set $E\subseteq X$, a strict total order $\succ$ over $X$ and a reference-dependent attention function $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that for all $A\in\mathcal{X}\setminus\phi$ and $x\in A$:
@@ 3 Characterization
In the present section I characterize the RAM-UR by using three testable axioms. These axioms restrict how the stochastic choice function operates across different menus.
Existence of a Dominant Alternative (EDA): For each $A\in\mathcal{X}\setminus\phi$: $p(x,A)=0$ for some $x\in A^{*}$ implies $p(y,\{x,y\})=1$ for some $y\in A\setminus\{x\}$.
@@
Even though the axioms restrict choice only where deterministic choice or non-choice is observed, it is not necessary to observe certain choice to identify the set of references. It is also not necessary to observe all menus to make such an identification. The following example illustrates that:
Example 5
Suppose we have four alternatives $X=\{a,b,c,d\}$ and the following observed choice probabilities:
$p(a,\{a,b,c\})=0.4\hskip 14.22636ptp(b,\{a,b,c\})=0.3\hskip 14.22636ptp(c,\{a,b,c\})=0.3$
$p(a,\{a,c,d\})=0.6\hskip 14.22636ptp(c,\{a,c,d\})=0.4\hskip 14.22636ptp(d,\{a,c,d\})=0$
$p(a,\{a,b,d\})=0.4\hskip 14.22636ptp(b,\{a,b,d\})=0.3\hskip 14.22636ptp(d,\{a,b,d\})=0.2$
$p(b,\{b,c,d\})=0.5\hskip 14.22636ptp(c,\{b,c,d\})=0.5\hskip 14.22636ptp(d,\{b,c,d\})=0$
Here we observe data only from menus of size three, and we do not observe any deterministic choice.
Note that in the above example, we can identify alternative $c$ as the only reference, as well as the partial order without the need for the complete set of menus. In fact, in any setting with a large number of alternatives and relatively few references, we can make full identification with very few data points, and we will not need to observe a lot of deterministic choices in those menus.
@@ 4 Random Attention Model with Unobserved References: Independent Random Attention
In this section I assume that the attention function in the RAM-UR model is of an independent random type in the sense that there exists an attention parameter $\gamma:X\rightarrow(0,1]$ such that the attention function takes the form: $\mu_{IRA}(D,A)=\prod_{x\in D\subseteq A}\gamma(x)\prod_{y\in A\setminus D}(1-\gamma(y))$ for all $A\in\mathcal{X}\setminus\phi$. Here attention privilege manifests in the form of alternatives with the attention parameter to be one, which is the key difference between this form and the attention function of the IRA form introduced by Manzini and Mariotti (2014) (hereafter referred to as the IRA model). The IRA restricts the attention parameter to be strictly between zero and one while we do not impose such a restriction. This makes the IRA model to be nested in the RAM-UR-IRA.
@@
A stochastic choice function $p$ is a random attention model with unobservable references and independent random attention (RAM-UR-IRA) if there exists a strict total order $\succ$ over $X$ and an attention parameter $\gamma:X\rightarrow(0,1]$ such that for all $A\in\mathcal{X}\setminus\phi$ and $x\in A$:
@@
Imposing the above additional structure significantly changes how the probabilities can be defined in the RAM-UR-IRA model. Even though it is a special case of RAM-UR and nests [^11] By nesting the IRA (rational) I mean that a special case of RAM-UR-IRA: the one where $E=\phi$ ($E=X$) will also be an IRA (rational) model. both the IRA as well as classical rationality, it is mathematically a small departure from the IRA model: It allows $\gamma$ to take value one for some alternatives. In what follows, I provide a characterization of this model by using axioms similar to the characterizations of both the IRA and the RAM-UR and discuss the identification results.
@@ 4.1 Characterization
R-Asymmetry (R-ASYM): For all $A\in\mathcal{X}\setminus\phi$ and $a,b\in A$ with $p(a,A)>0$, $p(b,A)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}\neq 1$ implies $\frac{p(b,A\setminus\{a\})}{p(b,A)}=1$.
@@
R-Independence (R-IND): (1) For all $A,B\in\mathcal{X}\setminus\phi$ and $a,b\in A\cap B$ with $p(a,A)>0,p(a,B)>0$: $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{p(a,B\setminus\{b\})}{p(a,B)}$.
(2) For all $A,B\in\mathcal{X}\setminus\phi$ and $b\in A\cap B$ with $p(a^{*},A)>0,p(a^{*},B)>0$: $\frac{p(a^{*},A\setminus\{b\})}{p(a^{*},A)}=\frac{p(a^{*},B\setminus\{b\})}{p(a^{*},B)}$.
@@
Existence of a Dominant Alternative* (EDA*): For all $A\in\mathcal{X}\setminus\phi$: $p(a^{*},A)=0$, implies $p(a^{*},\{x\})=0$ for some $x\in A$.
@@
Regularity (REG): For all $A,B\in\mathcal{X}\setminus\phi$: $A\subset B$ implies $p(a,A)\geq p(a,B)$ for all $a\in A^{*}$.
@@ 5 Literature Overview
Models of stochastic choice ventured into modelling the variability observed in choices due to limited attention. Manzini and Mariotti (2014) was the seminal paper that filled this void. Since then there has been a substantial body of literature combining both limited attention and stochastic choice. A plethora of papers are closely related to Manzini and Mariotti (2014), for instance, Brady and Rehbeck (2016)’s feasibility model is a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility. Another one is Horan (2019) which looks at the IRA without observing the probability of default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and models a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. They discussed the nestedness properties among these special cases. They were unable to conduct the nestedness analysis with respect to the classical stochastic choice models because in their model, the primitives were menu-reference pairs, which placed it outside the realm of stochastic choice models.
The idea of reference-dependent choice has also been the focus of several papers. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^13] Masatlioglu and Ok (2005) provides references for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. They model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They study the existence of a linear order such that the choice from each menu-reference pair is at least as good as the reference point, though not necessarily the maximum [^14] The best alternative according to the preference ordering.. The RAM-UR does have this non-deteriorating property since references are attention privileged, so in any menu, an alternative worse than the reference is never chosen.
@@
We have assumed menu-independence with respect to the references. In Kovach and Suleymanov (2023), the input to the problem is menu-reference pairs so the question of menu independence was moot. However, it is possible to allow for menu-dependent unobservable references as a possible extension to the model in the current paper. This extension obviously has bite, i.e., we cannot explain any stochastic choice function through such a model. For example, any stochastic choice function that allows for alternative $x$ to be chosen with certainty in a menu where $y$ is present and allows alternative $y$ to be chosen with certainty in a menu where $x$ is present cannot be a reference-dependent RAM even with menu-dependent references.
The literature on marketing and consumer research focused on consideration sets and the impact of memory and attention on whether certain brands enter consumers’ consideration sets. Nedungadi (1990) shows that whether an alternative is recalled while forming consideration sets is not necessarily correlated with how highly that alternative is evaluated by the consumer, i.e. attention might be independent of preferences. Advertising often plays a crucial role in increasing an alternative’s possibility of consideration, by increasing its brand recall, or in the terminology of our paper, making it a reference. It is also possible that higher advertising, and thus, more attention, increases the probability of choice of another alternative. This is consistent with both the experiments of Huber et al. (1982) and Nedungadi (1990), and something which can be observed in the model described in this paper. If a reference is added to a menu, the probability of choice of a different alternative can increase. The model described in this paper also seems relevant for more important decision problems like university choice for high school students (Dawes and Brown (2005), Rosen et al. (2008)) and political choice (Wilson (2008)). In these contexts, a fixed set of reference alternatives also makes sense.
Finally, the presence of a default alternative in our model requires some justification. In models of limited attention, the default is used to model the possibility of choice avoidance, delaying choice, which is quite a well-established empirical observation (Dhar (1997)). Horan (2019) argued that it might not be possible to observe the probability of no choice from empirical choice data, and provided a characterization of the IRA by Manzini and Mariotti (2014) without the need for observing such probabilities. However, note that in this model, the default is a de facto reference, and as such, any menu where another reference alternative is present, the default is not relevant. A reference in any menu is the alternative to which any alternative which is considered is always evaluated against. The separation of consideration and evaluation elements of the choice means that the default then becomes the reference for salience. The probability of choosing to abstain is the probability that none of the alternatives will be salient enough to draw attention. If there are other references present, those that are salient enough to always draw attention, the default is redundant.
@@
I take the set $\bar{E}=\{x\in X|p(x,\{x\})=1\}$ to be the revealed set of reference alternatives. If This set is empty EDA ensures that for any $A\in\mathcal{X}\setminus\phi$, $p(a,A)>0$ for all $a\in A^{*}$. In this case, we can construct an attention function with $E=\bar{E}=\phi$ and an arbitrary preference. Note that with an empty set of references, the RAM-UR is just a RAM, so it suffices to show a RAM representation. For this, I take an arbitrary preference and construct an attention function as follows: For any alternative the probability that it is chosen is taken to be the sum of the value that the attention function takes on subsets where that alternative is the best. One way to do it is divide the probability of an arbitrary alternative $x$ in a menu $A$ to all sets in $\{D\subseteq A:x\succ y,\forall y\in D\setminus\{x\}\}$ equally ($\succ$ is the arbitrary preference). Note that since all alternatives (and default) are chosen with positive probability in all menus such a construction will give us a valid attention function. Given the constructed attention function, the observed choices are represented by a RAM with the attention function and the arbitrary preferences. Such a construction will be one possible RAM representation. We can have multiple representations if we change how we assign the probabilities. Thus, we have a representation when $\bar{E}$ is empty. For the rest of the proof I assume that $\bar{E}$ is non-empty.
@@
R-ASYM: If there exist $A\in\mathcal{X}\setminus\phi$ and $x,y\in A$ with $p(x,A)>0$ and $p(y,A)>0$, then either none of $x$ and $y$ are references, in which case the axiom is true because of the functional form, as proved in Manzini and Mariotti (2014), or exactly one of them is a reference, and the other is preferred to it. In the latter case, the antecedent is true for the reference, since the probability of choice of the less preferred alternative (the reference) will increase if the other is removed, but removing the reference will not affect the choice probability of the better alternative.
R-IND: If there exists $A,B\in\mathcal{X}\setminus\phi$, $a,b\in A\cap B$ with $p(a,A)>0$ and $p(a,B)>0$, the functional form is such that the R-IND will hold. Same is true if instead of $a$ we consider the default $a^{*}$. The proof is identical to the proof of the necessity of i-independence in Manzini and Mariotti (2014).
@@
EDA*: If $p(a^{*},A)=0$, for some $A\in\mathcal{X}\setminus\phi$ then there exists at least one alternative $x\in A$ such that $\gamma(x)=1$. Then $p(a^{*},\{x\})=0$ follows.
@@
I take the set $\bar{E}=\{x\in X:p(x,A)=1\hskip 1.42262pt\text{for some}\hskip 1.42262ptA\in\mathcal{X}\setminus\phi\}$ to be the revealed set of reference alternatives.
Again, this set could be empty. If this set is empty, then EDA* and NT ensure that $p(a,A)>0$ for all $A\in\mathcal{X}\setminus\phi$ and $a\in A^{*}$. Now the R-ASYM and the R-IND will be equivalent to the i-ASYM and i-IND by Manzini and Mariotti (2014), which gives us the required equivalence with the IRA model [^16] An IRA model is a RAM-UR-IRA model with $E=\phi$.. For the remainder of the proof, I assume non-emptyness of $\bar{E}$.
@@
Reference-Independence of Irrelevant Alternatives (RIIA): For all $A\in\mathcal{X}\setminus\phi$ and $a\in A$: $p(a,A)=0$ implies $p(b,A\setminus\{a\})=p(b,A)$ for all $b\in A\setminus\{a\}$.
Claim 1: If a stochastic choice function $p:X\times\mathcal{X}\setminus\phi\rightarrow[0,1]$ satisfies REG, then it satisfies RIIA.
@@
Expansion (EXP): For all $A,B\in\mathcal{X}\setminus\phi$, $x_{1}\in A,x_{2}\in B$: $p(x_{1},A)=p(x_{2},B)=1$ implies either $p(x_{1},A\cup B)=1$ or $p(x_{2},A\cup B)=1$.
Claim 2: If a stochastic choice function $p:X\times\mathcal{X}\setminus\phi\rightarrow[0,1]$ satisfies R-ASYM, R-IND and REG, then it satisfies EXP.
@@
$A\cap\bar{E}=\phi$ for some $A\in\mathcal{X}\setminus\phi$ implies $p(a,A)>0$ for all $a\in A$.